How would I solve this?
The volume of a rectangular prism is 144 in³. What is the volume of a pyramid which has the same height and a congruent base?

Answer by poliphob3.14(115)

Solution:The volume of a pyramid with the same base and height is 1/3 of the volume of prism.
V(pyr.)=1/3V(prism)=1/3(144)=48 cubic inches.
Done.